OpenStudy (anonymous):

Factor this polynomial: 15/4-x-x^2

OpenStudy (xapproachesinfinity):

is that \[\frac{15}{4-x-x^2}\]

OpenStudy (xapproachesinfinity):

the way you wrote is vague

OpenStudy (anonymous):

\[\frac{ 15 }{ 4 } -x-x^2\] is what I meant

OpenStudy (xapproachesinfinity):

to factor that need to get to ax^2+bx+c form

OpenStudy (xapproachesinfinity):

what way do you know of for factoring

OpenStudy (anonymous):

I know, I can do that.\[-x^2-x+\frac{ 15 }{ 4 }\]

OpenStudy (xapproachesinfinity):

rather i want to say what ways

OpenStudy (anonymous):

Throw some at me, I can't really name 'em all...

OpenStudy (xapproachesinfinity):

that won't help you that much it is still the same

OpenStudy (xapproachesinfinity):

the most used one is quadratic

OpenStudy (xapproachesinfinity):

formula

OpenStudy (xapproachesinfinity):

we can do this \[1/4(-4x^2-4x+15)\]

OpenStudy (anonymous):

I know it's the same, you said I needed to put it into ax^2+bx+c form, so I did. I know quadratic formula.

OpenStudy (xapproachesinfinity):

can you factor now using any other method you know of

OpenStudy (xapproachesinfinity):

that's not what i meant see my comment you just rearranged the terms in descending order

OpenStudy (xapproachesinfinity):

\[1/4(-2x ~~~~~~~)(2x~~~~~~~)\]

OpenStudy (xapproachesinfinity):

what do you think will be in the blank

OpenStudy (anonymous):

5 and 3?